Find the mean and standard deviation of four-digit uniformly distributed lottery numbers (0000 through 9795). (Round your answers to 2 decimal places.)

Please show the steps taken and the formulas used, such as how you would enter the formula into excel.
Answer #1

here this is discrete uniform distribution ; with parameter a=0000 and b=9795

hence mean =(a+b)/2 =(0000+9795)/2 =4897.50

and std deviation =\sqrt{\frac{(b-a+1)^{2}-1}{12}} =\sqrt{\frac{(9795-0+1)^{2}-1}{12}} =2827.86

( please revert fir any clarfiication required)